How long would it take to double your principle (Just any number) at an annual interest rate of 8% compounded continuously.
Answer by lwsshak3(11628)

How long would it take to double your principle (Just any number) at an annual interest rate of 8% compounded continuously.
***
Formula for continuous compounding: A=Pe^rt, P=initial investment, r=interest rate, t=years, A=amt after years.
For doubling the principal, A/P=2=e^rt
e^rt=2
take log of both sides
rt*lne=ln2
lne=1
t=ln2/r==ln2/.08=8.66
How long would it take to double your principle? approximately 9 years